# Disentangling Tabular Data towards Better One-Class Anomaly Detection

### v1

@@ Abstract
Tabular anomaly detection under the one-class classification setting poses a significant challenge, as it involves accurately conceptualizing “normal” derived exclusively from a single category to discern anomalies from normal data variations. Capturing the intrinsic correlation among attributes within normal samples presents one promising method for learning the concept. To do so, the most recent effort relies on a learnable mask strategy with a reconstruction task. However, this wisdom may suffer from the risk of producing uniform masks, i.e., essentially nothing is masked, leading to less effective correlation learning. To address this issue, we presume that attributes related to others in normal samples can be divided into two non-overlapping and correlated subsets, defined as CorrSets, to capture the intrinsic correlation effectively. Accordingly, we introduce an innovative method that disentangles CorrSets from normal tabular data. To our knowledge, this is a pioneering effort to apply the concept of disentanglement for one-class anomaly detection on tabular data. Extensive experiments on 20 tabular datasets show that our method substantially outperforms the state-of-the-art methods and leads to an average performance improvement of 6.1% on AUC-PR and 2.1% on AUC-ROC.
@@ 1 Introduction
Figure 1: Visualization of MCM (Yin et al. 2024)’s fifteen soft masks and our two attention maps on the Thyroid dataset, which comprises six attributes per data point. The presented masks and attention maps are averages derived from all normal training samples. In MCM’s masks, darker colors indicate greater masking ratios while a higher attention weight in our attention maps.
@@
Motivated by the aforementioned assumption, we propose a novel paradigm named Disent-AD, which efficiently disentangles CorrSets from normal tabular data and adequately captures the intrinsic correlation. For practical implementation, a two-head self-attention module is utilized to implicitly extract two distinct subsets in latent space. The attention maps depict the attributes of interest to the network. By learning two independent attention maps, the two attention heads focus on different regions of tabular data, enabling the implicit extraction of distinct attribute subsets. As shown in Figure 1(b), darker regions in attention maps indicate greater weights. Evidently, the attention module successfully concentrates on two non-overlapping attribute subsets. To ensure that disentangled subsets are correlated, a reconstruction task is performed to restore the original data by utilizing subset features extracted from either of the two attention heads. We train the attention module with the reconstruction task in an end-to-end fashion. During testing, samples with high reconstruction errors are detected as anomalies.
@@ 3.2 Overview
Figure 2: Illustration of our strategy. Red regions indicate anomaly attributes. The CorrSets of normal data are denoted as $s_{1}^{norm}$ and $s_{2}^{norm}$. By learning the intrinsic correlation, each subset can independently restore original data. For anomalies, anomaly attributes disrupt the correlation, causing the reconstruction of the original data to fail.
Our strategy for anomaly detection is illustrated in Figure 2. Considering a tabular dataset where each sample is composed of $M$ attributes: $x_{i}=\{a_{1},a_{2},\cdots,a_{M}\}$. We denote the CorrSets of normal data as $s^{norm}_{i,1}$ and $s^{norm}_{i,2}$, $s^{norm}_{i,1}\cup s^{norm}_{i,2}=x_{i}^{norm}$ if there are no noise attributes in $x_{i}$. The inherent correlation can be captured by training network $f_{\theta}$ to restore the original data with only $s^{norm}_{1}$ or $s^{norm}_{2}$, i.e. $f_{\theta}(s^{norm}_{i,1})\rightarrow x_{i}^{norm}$ and $f_{\theta}(s^{norm}_{i,2})\rightarrow x_{i}^{norm}$. For anomaly instances during the test stage, we also extract two subsets, $s_{1}^{anom}$ and $s_{2}^{anom}$. If anomaly attributes are presented in $s_{1}^{anom}$ or $s_{2}^{anom}$, normal attributes of anomalies can not be reconstructed well, as the internal relations of anomalies deviate from those of normal instances. There is a case where anomalous attributes are not in two subsets, but the model fails to restore them using normal attribute subsets, still producing large reconstruction errors. Noisy attributes are unavoidable in real-world scenarios; however, neither normal nor anomaly attributes reconstruct noise well, which barely affects the distinction between normal and abnormal samples. Accordingly, a criterion based on the reconstruction error for anomaly detection can be built by disentangling CorrSets from normal samples.
We present the framework of our method in Figure 3. Given one sample for illustrating our procedure, the encoder takes it as input and extracts its latent features. Then, a two-head self-attention module is utilized to disentangle tabular data for extracting features of two attribute subsets. Afterward, a decoder maps these features back to the input space for independently reconstructing the original data. Our model is trained in an end-to-end fashion and optimized by a disentangling loss and a reconstruction loss. During inference, the sum of the two subsets’ reconstruction errors serves as the anomaly score for each test sample.
Figure 3: Illustration of our framework. The model consists of three modules: an encoder, a two-head self-attention, and a decoder. After extracting features of original data by the encoder, a two-head self-attention is performed for implicit disentanglement to obtain features of two subsets. They are then fed into the decoder to reconstruct the input data, respectively. The network is trained under the constraint of a disentangling loss and a reconstruction loss.
@@ 3.3 Our Proposed Method
We apply the attention module to latent features extracted by the encoder to disentangle tabular data. To successfully adopt the attention module, we add an additional dimension to tabular data, that is, $x_{i}\in\mathbb{R}^{M\times 1}$, where $M$ still denotes the number of attributes, and each attribute owns $1$ feature channel. The encoder $f_{E}$ maps $x_{i}$ to latent features $z_{i}\in\mathbb{R}^{M\times C}=f_{E}\left(x_{i}\right)$, where $C$ refers to the channel number of each attribute. With the attention, $z_{i}$ yields queries, keys, and values, symbolized respectively as $q_{i}^{s_{h}}$, $k_{i}^{s_{h}}$, $v_{i}^{s_{h}}$ $\in\mathbb{R}^{M\times C}$, for each head $h\in\{1,2\}$. The computation of the two attention maps is formalized as follows:
@@
Figure 4: Averaged reconstruction error and disentangling loss of test normal/anomaly samples during training on the Thyroid dataset.
@@ 4.3 Implementation Details
In our network architecture, a three-layer Multilayer Perceptron (MLP) with LeakyReLU activation function forms the encoder, and the decoder is symmetrically designed to the encoder. For the adaption of the attention module to tabular data, we add one additional dimension to most datasets and perform a patch-splitting preprocessing method for the rest to facilitate disentanglement, mostly derived from image scenarios. We consider the data structures of these datasets to be too complicated to affect disentangling. The preprocessing method is performed by splitting the data into three patches with the size of $M/2$ and treating each patch as an attribute to compose new data to simplify the data structure. We present a detailed description of the preprocessing method in Supplementary A. For datasets applied with the preprocessing method, we set epochs to 200 and the channel number $C$ of latent features to 512 for efficient convergence, while epochs to 100 and $C$ to 128 for the rest. Due to the large variation in the number of samples between datasets, ranging from 129 to 299,285, we use different batch sizes for different datasets. The parameters of the network are optimized by Adam with a uniform learning rate of 1e-4 for all datasets. See Supplementary B for more implementation details.
@@ 4.4 Baseline Methods
We compare our method with eleven outstanding tabular anomaly detection methods. IForest (Liu, Ting, and Zhou 2008), LOF (Breunig et al. 2000), OCSVM (Schölkopf et al. 1999), and ECOD (Li et al. 2022) represent classical non-deep methods. The competing deep methods include DAGMM (Zong et al. 2018), DeepSVDD (Ruff et al. 2018), GOAD (Bergman and Hoshen 2020), NeuTraLAD (Qiu et al. 2021), ICL (Shenkar and Wolf 2022), SLAD (Xu et al. 2023b), and MCM (Yin et al. 2024). It is noted that the results of IForest, LOF, OCSVM, ECOD, and DAGMM are obtained from MCM. We implement DeepSVDD, GOAD, NeuTraLAD, ICL, and SLAD by using DeepOD (Xu et al. 2023a), an open-source Python library for deep learning-based anomaly detection. The implementation of MCM is based on their official open-source code.
@@ 4.5 Anomaly Detection Performance
4.6 Ablation Study
We select four datasets sourced from different application scenarios to conduct the ablation study: Thyroid (healthcare), Arrhythmia (healthcare), Glass (forensic), and Satellite (satellite image). The results are reported in Table 3.
@@ 4.6 Ablation Study
Constraining the two subsets to be distinct is necessary for learning the correlation within normal data: A substantial performance drop is observed in “Variant 3”, since the absence of the disentangling loss hinders the model from extracting two non-overlapping attribute subsets. This indicates that the disentangling loss is crucial in our method.
@@ Appendix D Statistics of Datasets
Arrhythmia 452 274 66
Breastw 683 9 239
Cardio 1,831 21 176
Census 299,285 500 18,568
Campaign 41,188 62 4,640
Cardiotocography 2,114 21 466
Fraud 284,807 29 492
Glass 214 9 9
Ionosphere 351 33 126
Mammography 11,183 6 260
NSL-KDD 148,517 122 77,054
Optdigits 5,216 64 150
Pima 768 8 268
Pendigits 6,870 16 156
Satellite 6,435 36 2,036
Satimage-2 5,803 36 71
Shuttle 49,097 9 3,511
Thyroid 3,772 6 93
Wbc 278 30 21
Wine 129 13 10
@@
Figure 7: Illustration of the patch-splitting preprocessing.
(Epoch, $C$) Satellite Satimage-2
(100, 128) 88.3±0.1/85.2±0.3 97.0±0.1/99.1±0.0
(100, 512) 89.0±0.2/86.4±0.4 97.1±0.0/99.1±0.0
(200, 128) 89.0±0.0/86.2±0.1 97.3±0.4/99.3±0.3
(200, 512) 89.1±0.0/86.8±0.1 98.3±0.1/99.9±0.0

### v2

@@ Abstract
Tabular anomaly detection under the one-class classification setting poses a significant challenge, as it involves accurately conceptualizing “normal” derived exclusively from a single category to discern anomalies from normal data variations. Capturing the intrinsic correlation among attributes within normal samples presents one promising method for learning the concept. To do so, the most recent effort relies on a learnable mask strategy with a reconstruction task. However, this wisdom may suffer from the risk of producing uniform masks, i.e., essentially nothing is masked, leading to less effective correlation learning. To address this issue, we presume that attributes related to others in normal samples can be divided into two non-overlapping and correlated subsets, defined as CorrSets, to capture the intrinsic correlation effectively. Accordingly, we introduce an innovative method that disentangles CorrSets from normal tabular data. To our knowledge, this is a pioneering effort to apply the concept of disentanglement for one-class anomaly detection on tabular data. Extensive experiments on 20 tabular datasets show that our method substantially outperforms the state-of-the-art methods and leads to an average performance improvement of 6.1% on AUC-PR and 2.1% on AUC-ROC. Codes are available at https://github.com/yjnanan/Disent-AD.
@@ 1 Introduction
Figure 1: Visualization of MCM (Yin et al. 2024)’s fifteen soft masks and our two attention maps on the Thyroid dataset (six attributes). The presented masks and attention maps are averages derived from all training samples. In MCM’s masks, darker colors indicate greater masking ratios while a higher attention weight in our attention maps.
@@
Motivated by the aforementioned assumption, we propose a novel paradigm named Disent-AD, which efficiently disentangles CorrSets from normal tabular data and adequately captures the intrinsic correlation. For practical implementation, a two-head self-attention module is utilized to implicitly extract two distinct subsets in latent space. The attention maps depict the attributes of interest to the network. By learning two independent attention maps, the two attention heads focus on different regions of tabular data, enabling the implicit extraction of distinct attribute subsets. As shown in Figure 1(b), darker regions in attention maps indicate greater weights. Evidently, the attention module successfully concentrates on two non-overlapping attribute subsets. To ensure that disentangled subsets are correlated, a reconstruction task is performed to restore the original data by utilizing subset features extracted from either of the two attention heads. Importantly, this reconstruction process is strategically employed to detect anomalies rather than aiming for a complete restoration of the original data. During testing, samples with high reconstruction errors are detected as anomalies.
@@ 3.2 Overview
Figure 2: Illustration of our strategy. Red regions indicate anomaly attributes.
Our strategy for anomaly detection is illustrated in Figure 2. Considering a tabular dataset where each sample is composed of $M$ attributes: $x_{i}=\{a_{1},a_{2},\cdots,a_{M}\}$. We denote the CorrSets of normal data as $s^{norm}_{i,1}$ and $s^{norm}_{i,2}$, $s^{norm}_{i,1}\cup s^{norm}_{i,2}=x_{i}^{norm}$ if there are no noise attributes in $x_{i}$. The inherent correlation can be captured by training network $f_{\theta}$ to restore the original data with only $s^{norm}_{1}$ or $s^{norm}_{2}$, i.e. $f_{\theta}(s^{norm}_{i,1})\rightarrow x_{i}^{norm}$ and $f_{\theta}(s^{norm}_{i,2})\rightarrow x_{i}^{norm}$. For anomaly instances during the test stage, we also extract two subsets, $s_{1}^{anom}$ and $s_{2}^{anom}$. If anomaly attributes are presented in $s_{1}^{anom}$ or $s_{2}^{anom}$, normal attributes of anomalies can not be reconstructed well, as the internal relations of anomalies deviate from those of normal instances. There is a case where anomalous attributes are not in two subsets, but the model fails to restore them using normal attribute subsets, still producing large reconstruction errors. Noisy attributes are unavoidable in real-world scenarios; however, neither normal nor anomaly attributes reconstruct noise well, which barely affects the distinction between normal and abnormal samples. Accordingly, a criterion based on the reconstruction error for anomaly detection can be built by disentangling CorrSets from normal samples. It is worth noting that our method leverages reconstruction to detect anomalies rather than fully restoring the original data.
The framework of our method is presented in Figure 3. Given one sample for illustrating our procedure, the encoder takes it as input and extracts its latent features. Then, a two-head self-attention module is utilized to disentangle tabular data for extracting features of two attribute subsets. Afterward, a decoder maps these features back to the input space for independently reconstructing the original data. Our model is trained in an end-to-end fashion and optimized by a disentangling loss and a reconstruction loss. During inference, the sum of the two subsets’ reconstruction errors serves as the anomaly score for each test sample.
Figure 3: Illustration of our framework. The model consists of three modules: an encoder, a two-head self-attention, and a decoder. The network is trained under the constraint of a disentangling loss and a reconstruction loss.
@@ 3.3 Our Proposed Method
The attention module is applied to latent features extracted by the encoder to disentangle tabular data. To successfully adopt the attention module, we convert tabular data from row vectors into column vectors, i.e., $x_{i}\in\mathbb{R}^{M\times 1}$, where $M$ still denotes the number of attributes, and each attribute owns $1$ feature channel. The encoder $f_{E}$ maps $x_{i}$ to latent features $z_{i}\in\mathbb{R}^{M\times C}=f_{E}\left(x_{i}\right)$, where $C$ refers to the channel number of each attribute. With the attention, $z_{i}$ yields queries, keys, and values, symbolized respectively as $q_{i}^{s_{h}}$, $k_{i}^{s_{h}}$, $v_{i}^{s_{h}}$ $\in\mathbb{R}^{M\times C}$, for each head $h\in\{1,2\}$. The computation of the two attention maps is formalized as follows:
@@
Figure 4: Averaged reconstruction error and disentangling loss of test samples during training on the Thyroid dataset.
@@ 4.3 Implementation Details
In our network architecture, a three-layer Multilayer Perceptron (MLP) with LeakyReLU activation function forms the encoder, and the decoder is symmetrically designed to the encoder. For the adaption of the attention module to tabular data, we transform row vectors into column vectors for most datasets and perform a patch-splitting preprocessing method for the rest to facilitate disentanglement, mostly derived from image scenarios. We consider the data structures of these datasets to be too complicated to affect disentangling. The preprocessing method is performed by splitting the data into three patches with the size of $M/2$ and treating each patch as an attribute to compose new data to simplify the data structure. We present a detailed description of the preprocessing method in Supplementary A. For datasets applied with the preprocessing method, we set epochs to 200 and the channel number $C$ of latent features to 512 for efficient convergence, while epochs to 100 and $C$ to 128 for the rest. Due to the large variation in the number of samples between datasets, ranging from 129 to 299,285, we use different batch sizes for different datasets. The parameters of the network are optimized by Adam with a uniform learning rate of 1e-4 for all datasets. See Supplementary B for more implementation details.
@@ 4.4 Baseline Methods
Our method is compared with eleven outstanding tabular anomaly detection methods. IForest (Liu, Ting, and Zhou 2008), LOF (Breunig et al. 2000), OCSVM (Schölkopf et al. 1999), and ECOD (Li et al. 2022) represent classical non-deep methods. The competing deep methods include DAGMM (Zong et al. 2018), DeepSVDD (Ruff et al. 2018), GOAD (Bergman and Hoshen 2020), NeuTraLAD (Qiu et al. 2021), ICL (Shenkar and Wolf 2022), SLAD (Xu et al. 2023b), and MCM (Yin et al. 2024). It is noted that the results of IForest, LOF, OCSVM, ECOD, and DAGMM are obtained from MCM. We implement DeepSVDD, GOAD, NeuTraLAD, ICL, and SLAD by using DeepOD (Xu et al. 2023a), an open-source Python library for deep learning-based anomaly detection. The implementation of MCM is based on their official open-source code, and original results reported in their paper are presented in Supplementary E.
@@ 4.5 Anomaly Detection Performance
4.6 Ablation Study
Four datasets sourced from different application scenarios are selected to conduct the ablation study: Thyroid (healthcare), Arrhythmia (healthcare), Glass (forensic), and Satellite (satellite image). The results are reported in Table 3.
@@ 4.6 Ablation Study
Constraining the two subsets to be distinct is necessary for learning the correlation within normal data: A substantial performance drop is observed in “Variant 3”, since the absence of the disentangling loss hinders the model from extracting two non-overlapping attribute subsets. This indicates that the disentangling loss is crucial in our method. We report the results of replacing our disentangling loss with MCM (Yin et al. 2024)’s diversity loss in Supplementary E.
@@ 5 Conclusion
Acknowledgments
The work was partially supported by the following: National Natural Science Foundation of China under No. 92370119, No. 62376113, and No. 62206225; The Alan Turing Institute (UK) through the project ’Turing-DSO Labs Singapore Collaboration’ (SDCfP2\100009).
@@ Appendix D Statistics of Datasets
Appendix E More Results
MCM’s original results reported in their paper are presented in Table 8. Although MCM’s results are slightly higher than our reproduced results, our method still substantially outperforms it.
Table 9 reports AUC-PR/AUC-ROC results of replacing our disentangling loss with MCM’s diversity loss on Thyroid and Satellite datasets. Evidently, our method still outperforms MCM, demonstrating the superiority of our disentangling strategy.
@@ Appendix E More Results
Arrhythmia 452 274 66 (14.6%)
Breastw 683 9 239 (34.9%)
Cardio 1,831 21 176 (9.6%)
Census 299,285 500 18,568 (6.2%)
Campaign 41,188 62 4,640 (11.2%)
Cardiotocography 2,114 21 466 (22.0%)
Fraud 284,807 29 492 (0.1%)
Glass 214 9 9 (4.2%)
Ionosphere 351 33 126 (35.8%)
Mammography 11,183 6 260 (2.3%)
NSL-KDD 148,517 122 77,054 (51.8%)
Optdigits 5,216 64 150 (2.8%)
Pima 768 8 268 (34.8%)
Pendigits 6,870 16 156 (2.2%)
Satellite 6,435 36 2,036 (31.6%)
Satimage-2 5,803 36 71 (1.2%)
Shuttle 49,097 9 3,511 (7.1%)
Thyroid 3,772 6 93 (2.4%)
Wbc 278 30 21 (7.5%)
Wine 129 13 10 (7.7%)
@@
Figure 7: Illustration of the patch-splitting preprocessing.
(Epoch, $C$) Satellite Satimage-2
(100, 128) 88.3±0.1/85.2±0.3 97.0±0.1/99.1±0.0
(100, 512) 89.0±0.2/86.4±0.4 97.1±0.0/99.1±0.0
(200, 128) 89.0±0.0/86.2±0.1 97.3±0.4/99.3±0.3
(200, 512) 89.1±0.0/86.8±0.1 98.3±0.1/99.9±0.0
Dataset MCM MCM Ours
Arrhythmia 58.3±0.3/78.9±0.5 61.0/81.1 63.6±0.3/80.3±0.2
Breastw 99.7±0.0/99.7±0.0 99.5/99.5 99.8±0.0/99.7±0.0
Cardio 78.5±0.8/92.9±0.6 84.8/96.0 82.7±2.8/95.3±0.8
Census 21.3±0.8/73.6±0.6 24.2/75.8 39.7±0.0/83.3±0.0
Campaign 59.0±1.8/88.0±1.0 60.4/89.0 46.9±1.0/75.9±1.7
Cardiot. 69.7±1.4/81.1±1.2 69.9/80.0 75.1±0.9/81.9±0.2
Fraud 50.5±2.2/94.0±0.1 51.4/93.5 60.5±1.6/95.4±0.2
Glass 15.3±0.6/72.1±1.6 19.0/72.2 58.9±6.0/92.7±0.3
Ionosphere 97.7±0.1/96.7±0.1 98.0/97.2 98.3±0.2/97.8±0.2
Mammo. 47.9±6.4/91.2±1.5 47.5/90.5 43.9±1.5/90.2±0.3
NSL-KDD 90.2±0.6/87.9±0.4 90.8/87.8 88.6±0.8/83.2±0.5
Optdigits 80.2±8.1/98.9±0.5 88.8/99.4 86.8±4.8/99.1±0.2
Pima 69.4±0.9/70.9±1.5 73.8/76.3 73.7±2.7/75.8±2.2
Pendigits 74.2±10.0/98.9±0.5 82.5/99.1 89.5±6.6/99.5±0.3
Satellite 83.6±0.0/78.7±0.0 85.3/79.6 89.1±0.0/86.8±0.1
Satimage-2 97.8±0.2/99.8±0.0 98.5/99.9 98.3±0.1/99.9±0.0
Shuttle 97.8±0.5/99.9±0.0 94.7/99.7 99.5±0.2/99.9±0.0
Thyroid 83.0±0.5/97.8±0.0 84.1/98.0 89.4±0.7/98.9±0.0
Wbc 86.1±1.2/97.3±0.5 88.8/98.1 91.0±0.5/98.5±0.1
Wine 92.2±0.4/95.4±0.9 93.3/95.3 99.9±0.0/100.0±0.0
Average 72.6±1.8/89.6±0.5 74.8/90.4 78.7±1.5/91.7±0.3
Table 8: AUC-PR/AUC-ROC results of MCM and our method, †refers to our reproduced results, * denotes original results reported in their paper.
Method Thyroid Satellite
MCM 83.0/97.8 83.6/78.7
Ours (diversity loss) 88.3/98.6 88.9/86.2
Ours 89.4/98.9 89.1/86.8
Table 9: AUC-PR/AUC-ROC results of replacing our disentangling loss with MCM’s diversity loss.